OpenStudy (anonymous):

Let g(x)=2x^2+2x+1. Find the value of c between 1 and 3 such that the average rate of change of g(x) from x=1 to x=3 is equal to the instantaneous rate of g(x) at x=c.

OpenStudy (anonymous):

do just plug in 1 and 3 and do the rate of change and then plug in x and c and do the same?

myininaya (myininaya):

so do this \[g'(c)=\frac{g(3)-g(1)}{3-1}\]

OpenStudy (anonymous):

ok i get 10 doing that

myininaya (myininaya):

Have you plugged into the formula I gave you?

myininaya (myininaya):

find g' then find g'(c) and set that equal to the slope of the secant line from x=3 to x=1

OpenStudy (anonymous):

you're losing me lol

myininaya (myininaya):

do you know hot to find g'?

OpenStudy (anonymous):

4x + 2

OpenStudy (anonymous):

and then plug in 1 and 3?

myininaya (myininaya):

\[4c+2=\frac{g(3)-g(1)}{3-1}\]

OpenStudy (anonymous):

ok so the answer is 2

OpenStudy (anonymous):

thanks

myininaya (myininaya):

\[4c+2=\frac{2(3)^2+2(3)+1-[2(1)^2+2(1)+1]}{2}\]

OpenStudy (anonymous):

ok i have another similar prob that i need help with... ill post it on the main thing